# Changing the ranking in eigenvector centrality of a weighted graph by small perturbations

### v1

@@ 8. Conclusions
In this paper we have presented a methodology for finding a minimal perturbation to the adjacency matrix of a weighted (sparse) graph that causes the leading entries of the Perron eigenvector of the perturbed matrix to coalesce. The perturbation is constrained so as to preserve the sparsity of the original matrix. One of the motivations for this work was the question of the reliability and robustness of eigenvector centrality: if a small perturbation causes coalescence of the leading entries of the Perron eigenvector, this centrality measure cannot be considered reliable in the presence of uncertainties or noise, especially when the leading entries of the perturbed eigenvector differ from the original ones. Our methodology easily allows to incorporate additional constraints; for example, only a subset of the graph’s edge weights may be allowed to change. This feature can be used to identify which edges play an important role in influencing the ranking of nodes in the graph. The proposed algorithm appears to perform well on graphs arising from realistic applications, and is fairly efficient in spite of its complexity.

### v2

@@ 8. Conclusions
In this paper we have presented a methodology for finding a minimal perturbation to the adjacency matrix of a weighted (sparse) graph that causes the leading entries of the Perron eigenvector of the perturbed matrix to coalesce. The perturbation is constrained so as to preserve the sparsity of the original matrix. One of the motivations for this work was the question of the reliability and robustness of eigenvector centrality: if a small perturbation causes coalescence of the leading entries of the Perron eigenvector, this centrality measure cannot be considered reliable in the presence of uncertainties or noise, especially when the leading entries of the perturbed eigenvector differ from the original ones and the size of the perturbation is within the uncertainty limits. Our methodology easily allows to incorporate additional constraints; for example, only a subset of the graph’s edge weights may be allowed to change. This feature can be used to identify which edges play an important role in influencing the ranking of nodes in the graph. The proposed algorithm appears to perform favorably when compared with standard constrained optimization methods on graphs arising from realistic applications, and is fairly efficient in spite of its complexity.